OpenStudy (anonymous):

Find the general solution for each of the following linear differential equations. y′ +y=2 I have no clue what to do?

OpenStudy (anonymous):

this is a linear first order differential equation

OpenStudy (anonymous):

y'=2-y but y'=dy/dx therefore dy/(y-2)=-dx integrate it to get the answer

OpenStudy (anonymous):

a general solution will have a unknown constant let it be c

OpenStudy (anonymous):

So I just integrate dy/(y-2)=-dx ?

OpenStudy (anonymous):

yes thats it

OpenStudy (anonymous):

I got this y=2+Cexp(−x)from the answers does that seem right?

OpenStudy (anonymous):

that is right

OpenStudy (anonymous):

Thanks I appreciate the help!